# entity resolution AL active learning

### v2

@@ 4. Reusing of solved ER tasks
In this section, we introduce our novel method MoRER that takes advantage of previously solved ER problems, denoted as $\mathcal{P_{I}}$, along with their corresponding classification models. We initially give an overview and then describe the steps in detail.
@@ 4.1. Overview of MoRER
The similarities can be computed using appropriate similarity functions tailored to the relevant attributes of the records being compared. A further strategy to leverage data sources with different attributes is the use of pre-trained language models. The models can generate embeddings that encompass all attributes, functioning similarly to a bag-of-words approach. Using this strategy, an entity resolution approach can calculate embedding similarities for certain or all attributes.
Fig. 3 shows the workflow to efficiently create a model repository to solve new ER tasks of $\mathcal{P_{I}}$. We first conduct distribution analysis to determine whether the similarity distributions of features $f_{1},\ldots,f_{t}$ from the ER problems are similar. We then aggregate the results across the various features to compute an average distribution similarity score, $sim_{p}$, for each $p_{k,l}\in P_{I}$.
From the pairwise comparison of ER problems, we derive the ER problem similarity graph $G_{P}$. The ER problems represent the vertices, and the similarities between the ER problems are the weighted edges. We cluster the graph $G_{P}$ using the Leiden algorithm (Traag et al., 2019). In pre-experiments, we also tested different algorithms such as Girvan-Newmann (Girvan and Newman, 2002) and label propagation (Raghavan et al., 2007), leading to similar results.
@@
The result of the AL procedure for a cluster $C^{i}_{P}$ regarding the allocated budget is a labeled set of similarity feature vectors. We build a classification model with the obtained training data that we store for a certain cluster $C^{i}$. Moreover, we maintain the similarity feature vectors $P_{C^{i}}$ of the training data for each cluster to calculate similarities to new ER tasks.
Solving new ER problems $p_{x,z}\in\mathcal{P_{U}}$ requires selecting an appropriate model for each unsolved ER problem. We propose several strategies for this selection process. The first strategy, $sel_{\text{base}}$, identifies the most similar cluster $C^{i}$ and applies its associated model $M_{C^{i}}$. Specifically, the ER problem $p_{x,z}$ is compared with the set of similarity feature vectors $P_{C^{i}}$ from each cluster $C^{i}\in\mathcal{C_{P}}$. The cluster with the highest similarity score $sim_{p}$ is selected, and its corresponding model $M_{C^{i}}$ is used to classify the similarity vectors of $p_{x,z}$ into matches and non-matches.
@@
In the following section Sect. 4.2, we describe the method to compute similarities between ER problems. In detail, we provide a brief explanation of the distances used to compare two similarity distributions based on two features.
@@ 4.2. Similarity Distribution Analysis
Initially, we compare pairwise the existing ER tasks $p_{k,l}\in\mathcal{P_{I}}$ of the data sources being initially integrated. To simplify the comparison, we independently evaluate the distributions of each similarity feature $f_{1},\ldots,f_{t}$. The application of univariate distribution tests is less complex and computationally expensive compared to multivariate tests. Moreover, we assume that the feature spaces are of the same size. However, for highly heterogeneous data sources, we recommend generating record embeddings based on the available attributes for each data source and calculating similarities between these embeddings covering multiple attributes.
We compute a distribution similarity $sim_{dfi}$ for each feature pair and aggregate them into a weighted average similarity $sim_{p}$ using the standard deviation of each feature from the ER problems.
@@
Kolmogorow-Smirnow-Test. The Kolmogorov-Smirnov test (Hollander et al., 2013) measures the maximum distance between the cumulative distribution functions $CDF^{f}_{k,l}$ and $CDF^{f}_{m,n}$ considering a certain feature $f$ of the ER tasks $p_{k,l}$ and $p_{m,n}$. The test statistic is defined by the supremum of the absolute difference between the two cumulative distribution functions given by Eq. 1.
@@
Wasserstein distance. The Wasserstein distance (Villani, 2009) measures the cost to transform a distribution $Q$ into a distribution $W$. In our case, we consider the distributions $d^{f}_{k,l}$ and $d^{f}_{m,n}$ given by two column vectors for a feature $f$ of the entity resolution problems $p_{k,l}$, resp. $p_{m,n}$. Using these column vectors, the corresponding cumulative distribution functions $CDF^{f}_{k,l}$ and $CDF^{f}_{m,n}$ are calculated and adapted to the same size if the column-wise feature vectors have different dimensions. The Wasserstein distance is the sum of the absolute differences between corresponding elements of $CDF^{f}_{k,l}$ and $CDF^{f}_{m,n}$, illustrated in Eq. 2.
@@
Population stability index. The population stability index (Siddiqi, 2005) quantifies covariate shifts between two univariate distributions $d^{f}_{k,l}$ and $d^{f}_{m,n}$. Each distribution is split into bins $B_{k,l}$ and $B_{m,n}$ where the proportion $prop(b^{i}_{k,l})$ resp. $prop(b^{i}_{m,n})$ of similarities for the bins is determined. The index is defined as follows, where 100 is a commonly used number of bins $|B|$:
@@
The resulting distances of the features between two ER problems $p_{k,l}$ and $p_{m,n}$ are transformed into similarities and averaged as $sim_{p}$. Due to the different importance of similarity features, we weigh the average by utilizing the standard deviation of a similarity feature vector to consider the discriminative power of these features.
@@ 4.4. Model Generation
We consider active learning approaches and supervised ML methods for generating the corresponding model for each cluster. For the active learning approaches, we distribute a given budget $b_{tot}$ to the clusters.
Budget distribution
For each resulting cluster $C^{i}\in\mathcal{C_{P}}$, we generate a classification model $M_{C^{i}}$ to classify ER problems that are similar to those of cluster $C^{i}$. We limit the number of record pairs to build the training data by $b_{tot}$. We assign a minimum number of labels $b_{min}$ to each cluster to guarantee sufficient training data. To guarantee that the total budget is sufficient, we distinguish between singleton and non-singleton clusters $\mathcal{C_{P}}_{|s}$, respectively. $\mathcal{C_{P}}_{|ns}$. If Eq. 4 holds, we have to merge the clusters of $\mathcal{C_{P}}_{|s}$ into clusters $C^{i}\in\mathcal{C_{P}}_{|ns}$ since the total budget is not sufficient for all clusters.
@@
(6) $ratio(\mathcal{C_{P}}_{|ns})=\frac{|\mathcal{C_{P}}_{|ns}|}{|\mathcal{P}|}$
(7) $ratio(\mathcal{C_{P}}_{|s})=\frac{|\mathcal{C_{P}}_{|s}|}{|\mathcal{P}|}$
@@
Training data selection
A cluster $C^{i}$ consists of multiple ER problems that are similar to each other. We assume that one model can classify similar ER problems. Therefore, we train one classifier for each cluster. The model generation requires training data consisting of feature vectors labeled as matches and non-matches. In our method, we integrate two AL methods, Almser (see Sect. 3) and the uncertainty AL method from Mozafari et al. (Mozafari et al., 2014). The main idea of an AL method is to select iteratively a certain number batchbatch of unlabeled similarity feature vectors $w$ that are informative to extend the current training data $\mathbf{T}$. The iterative process terminates if a stop criterion, such as a total labeling budget, is reached. Note that other AL methods can be integrated. We have chosen Almser since it is the only method focusing on multi-source ER. Furthermore, the uncertainty method (Mozafari et al., 2014) scales for large ER problems and simultaneously achieves qualitative results.
@@
Almser: We use the original implementation [^1] https://github.com/wbsg-uni-mannheim/ALMSER-GB and extend it to support batch processing. As input, we only consider the induced subgraph regarding the ER tasks of a cluster.
@@ 4.5. Solving New Entity Resolution Problems
(14) $b_{\text{new}}=b_{\text{tot}}\cdot cov(C)\cdot\frac{|\{w\mid w\in T\cap C^{\text{prev}}\}|}{b_{\text{tot}}}$
@@ 5. Evaluation
In this section, we evaluate our method MoRER using three datasets, considering the multi-source AL method Almser (Primpeli and Bizer, 2021), the transfer learning method TransER (Kirielle et al., 2022) focusing on relational ER tasks, the self-supervised approach Sudowoodo (Wang et al., 2023), the unsupervised approach MultiEM, the SOTA supervised ER approach Ditto (Li et al., 2020), and the small language model-based method AnyMatch (Zhang et al., 2025). Moreover, we analyze the components of MoRER such as the various distribution tests, the amount of initially solved ER problems, and the impact of retraining.
@@ 5.1. Datasets
Our method focuses on multi-source ER approaches, where the availability of datasets is limited. We consider three datasets being widely used in this domain (Saeedi et al., 2018; Zeng et al., 2024; Primpeli and Bizer, 2021; de Meer Pardo et al., 2025) as summarized in Table 2.
@@ 5.2. Experimental Setup
We implemented MoRER in Python 3.12 using scikit-learn 1.5.1 and networkx 3.3. We modified the existing Almser implementation as available on GitHub and reimplemented the uncertainty active learning method from Mozafari et. al. (Mozafari et al., 2014), which we called bootstrap. The approach determines the uncertainty using k models. We set k=100 achieving qualitative results following previous studies (Christen et al., 2019; Mozafari et al., 2014).
@@
Similarity Feature Vectors: The main focus of our method is the classification and, consequently, the model generation and selection. Therefore, we pre-calculate the similarity vectors of the ER problems. We use the same attribute comparisons and blocking keys for the Dexter dataset as in previous studies (Lerm et al., 2021; Saeedi et al., 2021). For the WDC-computer and Music datasets, we utilize the feature vectors [^2] http://data.dws.informatik.uni-mannheim.de/benchmarkmatchingtasks/almser_gen_data/ used in the previous work presenting Almser  (Primpeli and Bizer, 2021). The authors use various string similarity functions for textual data and normalized absolute differences for numerical values.
@@
Method Parameters: We use the following configuration summarized in Table 3, where the values written in bold are default values.
@@
Figure 5. Linkage quality comparison of MoRER to Almser standalone, Sudowoodo, AnyMatch, TransER, and Ditto.
Figure 6. Runtime comparison of MoRER to Almser standalone, TransER, AnyMatch, Sudowoodo, and Ditto.
Compared methods
Due to the limitations of existing methods for reusing ML models, we focused on label-efficient techniques like active learning and transfer learning. Our method is compared to Almser  (Primpeli and Bizer, 2021), and TransER  (Kirielle et al., 2022). All methods utilized identical similarity feature vectors, with a consistent division between solved and unsolved ER tasks to ensure comparability.
We also compare our method against language model-based approaches, which have shown superior performance over traditional methods for processing textual data. To generate record embeddings, we utilize the same set of attributes originally employed to create the similarity feature vectors. We consider MultiEM  (Zeng et al., 2024) as a representative of unsupervised approaches, Sudowoodo  (Wang et al., 2023) as a self-supervised learning method, Ditto  (Li et al., 2020) as a state-of-the-art supervised approach, and AnyMatch  (Zhang et al., 2025) using small language models.
We utilize the same record pairs for the training and test sets for each method. We partition the training and test data sets by the source pairs of the corresponding record pairs to generate ER problems for $\mathcal{P_{I}}$ and $\mathcal{P_{U}}$. Unlike AL methods, such as Almser, TransER, and Ditto are not designed to select informative links. Instead, we treat all initial ER tasks $\mathcal{P_{I}}$ among the integrated data sources as training data and 50% of the similarity feature vectors for training. This consideration results in the following training data sizes of all pairs for the Dexter, WDC-computer, and Music datasets: 68,339, 231,521, and 46,049 record pairs, respectively. These training sizes reflect the volume of labeled record pairs used by TransER and Ditto for their classification tasks. For Sudowoodo, we use the semi-supervised approach with the same labeling budget as our method MoRER and Almser.

### v3

@@ 4. Reusing of Solved ER Tasks
In this section, we introduce our novel method MoRER that takes advantage of previously solved ER problems, denoted as $\mathcal{P_{I}}$, along with their corresponding classification models. We initially give an overview and then describe the steps of MoRER in detail.
@@ 4.1. Overview of MoRER
The similarities can be computed using appropriate similarity functions tailored to the relevant attributes of the records being compared (Christen, 2012). A further strategy to leverage data sources with different attributes is the use of pre-trained language models. The models can generate embeddings that encompass all attributes, functioning similarly to a bag-of-words approach. Using this strategy, an ER can calculate embedding similarities for certain or all attributes.
Figure 3 shows the workflow to efficiently create a model repository to solve new ER tasks of $\mathcal{P_{I}}$. We first conduct distribution analysis to determine whether the similarity distributions of features $f_{1},\ldots,f_{t}$ from the ER problems are similar considering univariate and multivariate analysis tests. We calculate an similarity score, $sim_{p}$, for each $p_{k,l}\in\mathcal{P_{I}}$.
From the pairwise comparison of ER problems, we derive the ER problem similarity graph $G_{P}$. The ER problems represent the vertices, and the similarities between the ER problems are the weighted edges. We cluster the graph $G_{P}$ using the Leiden algorithm (Traag et al., 2019). In pre-experiments, we also evaluated other algorithms such as Girvan-Newmann (Girvan and Newman, 2002) and label propagation (Raghavan et al., 2007), leading to similar results. As shown in Fig. 3, we calculate pairwise the similarities between $P_{1,2},P_{1,3}\ldots P_{m,n}$ and compute clusters based based on the graph $G_{P}$ resulting in the clusters $C_{1},C_{2}$, and $C_{3}$.
@@
The result of the AL procedure for a cluster $C^{i}_{P}$ regarding the allocated budget is a labeled set of similarity feature vectors. We build a classification model with the obtained training data that we store for a certain cluster $C^{i}$. Moreover, we maintain the similarity feature vectors $P_{C^{i}}$ of the training data for each cluster to calculate similarities to new ER tasks. Considering the example in Fig. 3, we generate the corresponding models $M_{C1},M_{C2}$ and $M_{C3}$ for the three clusters. Each cluster is represented by the training data that has been used for the corresponding similarity calculations.
Solving new ER problems $p_{x,z}\in\mathcal{P_{U}}$ requires selecting an appropriate model for each unsolved ER problem. We propose several strategies for this selection process. The first strategy, $sel_{\text{base}}$, identifies the most similar cluster $C^{i}$ and applies its associated model $M_{C^{i}}$. Specifically, the ER problem $p_{x,z}$ is compared with the set of similarity feature vectors $P_{C^{i}}$ from each cluster $C^{i}\in\mathcal{C_{P}}$. The cluster with the highest similarity score $sim_{p}$ is selected, and its corresponding model $M_{C^{i}}$ is used to classify the similarity vectors of $p_{x,z}$ into matches and non-matches. In the example, we calculate the similarities of the new ER problem $P_{5,6}$ to the cluster representatives $P_{C^{1}},P_{C^{2}}$ and $P_{C^{3}}$.
@@
In the following section, we describe the method to compute similarities between ER problems. In detail, we provide a brief explanation of the distances used to compare two similarity distributions based on two features.
@@ 4.2. Similarity Distribution Analysis
Initially, we compare pairwise the existing ER tasks $p_{k,l}\in\mathcal{P_{I}}$ of the data sources being initially integrated. We consider univariate and multivariate statistical analysis. For the univariate tests, we independently evaluate the distributions of each similarity feature $f_{1},\ldots,f_{t}$. The application of univariate distribution tests is less complex and computationally expensive compared to multivariate tests. Moreover, we assume that the feature spaces are of the same size. However, for highly heterogeneous data sources, we recommend generating record embeddings based on the available attributes for each data source and calculating similarities between these embeddings covering multiple attributes.
For the univariate tests, we calculate a distribution similarity $sim_{dfi}$ for each feature pair and aggregate them into a weighted average similarity $sim_{p}$ using the standard deviation of each feature from the ER problems.
@@
Kolmogorow-Smirnow-Test. This test (Hollander et al., 2013) measures the maximum distance between the cumulative distribution functions $CDF^{f}_{k,l}$ and $CDF^{f}_{m,n}$ considering a certain feature $f$ of the ER tasks $p_{k,l}$ and $p_{m,n}$. The test statistic is defined by the supremum of the absolute difference between the two cumulative distribution functions given by Eq. 1.
@@
Wasserstein distance. This distance (Villani, 2009) measures the cost to transform a distribution $Q$ into a distribution $W$. In our case, we consider the distributions $d^{f}_{k,l}$ and $d^{f}_{m,n}$ given by two column vectors for a feature $f$ of the entity resolution problems $p_{k,l}$, resp. $p_{m,n}$. Using these column vectors, the corresponding cumulative distribution functions $CDF^{f}_{k,l}$ and $CDF^{f}_{m,n}$ are calculated and adapted to the same size if the column-wise feature vectors have different dimensions. The Wasserstein distance is the sum of the absolute differences between corresponding elements of $CDF^{f}_{k,l}$ and $CDF^{f}_{m,n}$, illustrated in Eq. 2.
@@
Population stability index. This index (Siddiqi, 2005) quantifies covariate shifts between two univariate distributions $d^{f}_{k,l}$ and $d^{f}_{m,n}$. Each distribution is split into bins $B_{k,l}$ and $B_{m,n}$ where the proportion $prop(b^{i}_{k,l})$ resp. $prop(b^{i}_{m,n})$ of similarities for the bins is determined. The index is defined as follows, where 100 is a commonly used number of bins $|B|$:
@@
The resulting univariate distances of the features between two ER problems $p_{k,l}$ and $p_{m,n}$ are transformed into similarities and averaged as $sim_{p}$. Due to the different importance of similarity features, we weigh the average by utilizing the standard deviation of a similarity feature vector to consider the discriminative power of these features.
Classifier two-sample test. In addition to univariate distance measures, we employ a classifier two-sample test for multivariate distribution analysis (Lopez-Paz and Oquab, 2016). The core idea is to train a classification model to distinguish between two datasets. If the distributions are similar, the classifier’s accuracy will be low, indicating that it struggles to differentiate between samples from the two datasets. We use the F1 score (Christen et al., 2023) to address the skewness according to the size of ER tasks. We define $sim_{p}$ as the inverse F1 score.
@@ 4.4. Model Generation
We consider AL approaches and supervised ML methods for generating the corresponding model for each cluster. For the AL approaches, we distribute a given budget $b_{tot}$ to the clusters.
Budget distribution. For each resulting cluster $C^{i}\in\mathcal{C_{P}}$, we generate a classification model $M_{C^{i}}$ to classify ER problems that are similar to those of cluster $C^{i}$. We limit the number of record pairs to build the training data by $b_{tot}$. We assign a minimum number of labels $b_{min}$ to each cluster to guarantee sufficient training data. To guarantee that the total budget is sufficient, we distinguish between singleton and non-singleton clusters $\mathcal{C_{P}}_{|s}$, respectively. $\mathcal{C_{P}}_{|ns}$. If Eq. 4 holds, we have to merge the clusters of $\mathcal{C_{P}}_{|s}$ into clusters $C^{i}\in\mathcal{C_{P}}_{|ns}$ since the total budget is not sufficient for all clusters.
@@
(6) $ratio(\mathcal{C_{P}}_{|ns})=\nicefrac{{|\mathcal{C_{P}}_{|ns}|}}{{|\mathcal{P}|}}$
(7) $ratio(\mathcal{C_{P}}_{|s})=\nicefrac{{|\mathcal{C_{P}}_{|s}|}}{{|\mathcal{P}|}}$
@@
Training data selection. A cluster $C^{i}$ consists of multiple ER problems that are similar to each other. We assume that one model can classify similar ER problems. Therefore, we train one classifier for each cluster. The model generation requires training data consisting of feature vectors labeled as matches and non-matches. In our method, we integrate two AL methods, Almser (Primpeli and Bizer, 2021) (see Sect. 3) and the uncertainty AL method from Mozafari et al. (Mozafari et al., 2014). Note that other AL methods can be integrated. We have chosen Almser since it is the only method focusing on multi-source ER. Furthermore, the uncertainty method (Mozafari et al., 2014) scales for large ER problems and simultaneously achieves qualitative results.
@@
Almser. We use the original implementation [^1] https://github.com/wbsg-uni-mannheim/ALMSER-GB and extend it to support batch processing. As input, we only consider the induced subgraph regarding the ER tasks of a cluster.
@@ 4.5. Solving New Entity Resolution Problems
(14) $b_{\text{new}}=b_{\text{tot}}\cdot cov(C)\cdot\nicefrac{{|\{w\mid w\in T\cap C^{\text{prev}}\}|}}{{b_{\text{tot}}}}$
@@ 5. Evaluation
In this section, we evaluate our method MoRER using three datasets, considering the multi-source AL method Almser (Primpeli and Bizer, 2021), the transfer learning method TransER (Kirielle et al., 2022) focusing on relational ER tasks, the self-supervised approach Sudowoodo (Wang et al., 2023), the SOTA supervised ER approach Ditto (Li et al., 2020), and the small language model-based method AnyMatch (Zhang et al., 2025). Moreover, we analyze the components of MoRER such as the various distribution tests and the impact of retraining.
@@ 5.1. Datasets
Our method focuses on multi-source ER approaches, where the availability of datasets is limited. We consider three datasets being widely used in this domain (Saeedi et al., 2018; Zeng et al., 2024; Primpeli and Bizer, 2021; de Meer Pardo et al., 2025; Lerm et al., 2021) as summarized in Table 2.
@@ 5.2. Experimental Setup
We implemented MoRER in Python 3.12 using scikit-learn 1.5.1 and networkx 3.3. We modified the existing Almser implementation as available on GitHub and reimplemented the uncertainty active learning method from Mozafari et. al. (Mozafari et al., 2014), which we called Bootstrap. We set k=100 achieving qualitative results following previous studies (Christen et al., 2019; Mozafari et al., 2014). We evaluate the effectiveness by calculating precision, recall, and F1 score (Christen et al., 2023) according to the predicted matches across overall ER tasks for a test dataset.
@@
Similarity Feature Vectors. The main focus of our method is the classification and, consequently, the model generation and selection. Therefore, we pre-calculate the similarity vectors of the ER problems. We use the same attribute comparisons, string similarity functions, and blocking keys for the Dexter dataset as in previous studies (Lerm et al., 2021; Saeedi et al., 2021). For the WDC-computer and Music datasets, we utilize the feature vectors [^2] http://data.dws.informatik.uni-mannheim.de/benchmarkmatchingtasks/almser_gen_data/ used in the previous work presenting Almser  (Primpeli and Bizer, 2021). The authors use various string similarity functions for textual data and normalized differences for numerical values.
@@
Method Parameters. We use the following configuration summarized in Table 3, where the values written in bold are default values.
@@
Table 4. Linkage quality (Precision/Recall/F1 score) comparison of the budget limited methods MoRER in combination with an AL method to Almser standalone, Sudowoodo, and AnyMatch, and the supervised methods MoRER, TransER, and Ditto. We highlight the best F1 score in bold and underline the second best one for each budget and method type (AL vs supervised).
D B MoRER+Almser MoRER+BS Almser Sudowoodo AnyMatch B MoRER Ditto Unicorn TransER
D 1K 0.96/0.92/0.94 0.88/0.90/0.89 0.97/0.87/0.92 0.42/0.76/0.54 0.59/0.96/0.73 50% 0.94/0.92/0.93 0.60/0.99/0.75 0.53/0.96/0.68 0.95/0.90/0.92
1.5K 0.96/0.92/0.94 0.85/0.94/0.89 0.97/0.89/0.92 0.43/0.68/0.53 0.60/0.97/0.74 all 0.91/0.93/0.92 0.60/0.99/0.75 0.56/0.93/0.70 0.94/0.87/0.90
2K 0.96/0.90/0.93 0.88/0.93/0.90 0.97/0.90/0.94 0.45/0.85/0.58 0.59/0.98/0.73
W 1K 0.84/0.83/0.83 0.85/0.77/0.81 0.93/0.92/0.93 0.43/0.99/0.60 0.55/0.97/0.70 50% 0.94/0.93/0.94 0.92/0.95/0.93 0.77/0.92/0.84 0.79/0.86/0.82
1.5K 0.94/0.87/0.90 0.90/0.86/0.88 0.94/0.93/0.93 0.40/0.99/0.57 0.58/0.94/0.72 all 0.94/0.92/0.93 0.98/0.94/0.96 0.90/0.90/0.90 0.80/0.86/0.83
2K 0.95/0.94/0.94 0.79/0.84/0.84 0.95/0.93/0.94 0.47/0.90/0.61 0.61/0.98/0.75
M 1K 0.98/0.96/0.97 0.94/0.92/0.97 0.98/0.97/0.98 0.92/0.95/0.94 0.97/0.99/0.98 50% 0.99/0.96/0.97 0.99/0.99/0.99 0.98/0.98/0.98 0.96/0.93/0.94
1.5K 0.98/0.97/0.98 0.98/0.97/0.97 0.99/0.97/0.98 0.95/0.96/0.96 0.93/0.99/0.96 all 0.99/0.97/0.98 1.00/0.99/0.99 0.97/0.99/0.98 0.95/0.92/0.94
2K 0.99/0.96/0.98 0.96/0.96/0.97 0.99/0.97/0.98 0.93/0.96/0.95 0.96/0.99/0.97
Figure 5. Runtime comparison of MoRER with AL and the supervised version to Almser, Sudowoodo, AnyMatch, TransER, and Ditto. The shaded areas represent the runtime for the statistical analysis/clustering (striped) and the selection (dotted) of a model.
Compared methods. Due to the limitations of existing methods for reusing ML models, we focused on label-efficient techniques like active learning and transfer learning. Our method is compared to Almser  (Primpeli and Bizer, 2021), and TransER  (Kirielle et al., 2022). All methods utilized identical similarity feature vectors, with a consistent division between solved and unsolved ER tasks to ensure comparability.
We also compare our method against language model-based approaches, which have shown superior performance over traditional methods for processing textual data. To generate record embeddings, we utilize the same set of attributes originally employed to create the similarity vectors. We consider Sudowoodo  (Wang et al., 2023), Ditto  (Li et al., 2020), Unicorn  (Fan et al., 2024), and AnyMatch  (Zhang et al., 2025).
We utilize the same record pairs for the training and test sets for each method. We partition the training and test data sets by the source pairs of the corresponding record pairs to generate ER problems for $\mathcal{P_{I}}$ and $\mathcal{P_{U}}$. Unlike AL methods (such as Almser), TransER, Ditto, and Unicorn are not designed to select informative links. Instead, we treat all initial ER tasks $\mathcal{P_{I}}$ among the integrated data sources as training data and 50% of the similarity feature vectors for training. This consideration results in the following training data sizes of all pairs for the Dexter, WDC-computer, and Music datasets: 68,339, 231,521, and 46,049 record pairs, respectively. These training sizes reflect the volume of labeled record pairs used by TransER, Ditto, and Unicorn for their classification tasks. For Sudowoodo and AnyMatch, we use the semi-supervised approach with the same labeling budget as our method MoRER and Almser.